# MQFL-FHE: Multimodal Quantum Federated Learning Framework with Fully Homomorphic Encryption

### v3

@@ I Introduction
Federated Learning (FL) has emerged as a powerful paradigm for collaboratively training machine learning models across decentralized data sources ((McMahan et al., 2023)), allowing multiple clients to enhance model performance while preserving data privacy and complying with regulations such as the general data protection regulation (GDPR). FL alone is insufficient to fully protect data privacy, as the global model and updates shared between clients and the central server remain vulnerable to inference attacks. For instance, membership inference attacks use the model parameters or gradients to deduce whether a specific data sample was part of the training dataset ((Shokri et al., 2017)). Similarly, attribute inference attacks attempt to uncover sensitive attributes of individual data samples by exploiting correlations between model updates and data features ((Melis et al., 2019)). Fully homomorphic encryption (FHE) was proposed by (Gentry, 2009) to mitigate vulnerabilities by allowing computations on encrypted data without decryption, ensuring that sensitive information remains secure even during model training and aggregation within new gen FL. It works well against gradient leakage attacks, in which hackers try to recover private information by looking at the gradients that are sent back and forth during training ((Phong et al., 2018)). Adding FHE to FL adds a lot of extra work to computers, causing delays and less accurate models during the aggregation phase ((Zhang et al., 2023)). Consequently, while FHE protects data privacy, it can hinder the overall performance of FL systems, which is crucial for applications requiring high accuracy or real-time predictions. Despite the growing interest in multimodal FL, which aims to leverage various types of data from multiple clients, there exists a gap in research that addresses the unique challenges posed by FHE in this context ((Gong et al., 2024a)). Furthermore, the application of quantum computing (QC) to enhance FL frameworks has been largely unexplored. QC has the potential to accelerate specific computations, which could alleviate performance bottlenecks introduced by FHE ((Dutta et al., 2024)).
To mitigate the performance degradation that occurs during model aggregation, we propose a novel integration of QC with FHE in FL, taking into account the nature of the problem. Our method employs QC to counteract the model degradation resulting from FHE, allowing for an improved collection of model updates while protecting privacy, and for the first time, we present a novel multimodal quantum mixture of experts (MQMoE) model within the FL framework that employs FHE. We design this architecture to handle diverse multimodal data with MoE ((Yu et al., 2023a)) from various clients, enhancing representational and task-specific learning performance.
@@ II Related Works
QC in FL: (Innan et al., 2024a) introduced the federated quantum neural network (FedQNN), which integrates quantum machine learning (QML) with QFL principles ((Biamonte et al., 2017)). The experiments were conducted across genomics and healthcare datasets, laying the groundwork for investigations into quantum-enhanced learning methods in privacy-sensitive domains. This has been further advanced by efforts in QFL, which are creating new avenues in QML. In particular, studies like (Chen and Yoo, 2021; Chehimi and Saad, 2022) have emphasized using QFL in quantum systems and decentralized data frameworks. Applications in healthcare and other domains, using both quantum hardware and classical simulations, have shown promising results ((Lusnig et al., 2024; Bhatia and Bernal Neira, 2024; Innan et al., 2024b; Javeed et al., 2024; Bhatia et al., 2023; Kais et al., 2023)). Despite this progress, challenges related to resource allocation and efficient implementation persist, as noted by (Larasati et al., 2022; Ren et al., 2023; Gurung et al., 2023), highlighting ongoing efforts to address these issues.
Privacy preservation in FL: (Rahulamathavan et al., 2023) proposed FheFL, which incorporates a multi-key additive homomorphic encryption scheme to mitigate privacy risks associated with gradient sharing in FL. The experimental results confirmed that FheFL maintains model accuracy while providing robust security against potential data breaches, emphasizing the necessity of integrating FHE techniques into FL frameworks to enhance data privacy. FedSHE was created by (Pan et al., 2024). It is a useful FL scheme that uses adaptive segmented CKKS homomorphic encryption to keep gradients safe from leakage attacks. Their analysis of security parameters underscores the trade-offs between correctness, efficiency, and security in FL environments.
Multimodal data handling in FL: (Yu et al., 2023b) developed CreamFL, a framework designed for multimodal FL. This method enables clients with heterogeneous model architectures to contribute to a unified global model without sharing their private data. By implementing a contrastive representation ensemble strategy, CreamFL effectively addresses the challenges posed by modality discrepancies and task diversities. (Gong et al., 2024b) further advanced this area by introducing a multimodal vertical FL framework utilizing bivariate Taylor series expansion for encryption, thus eliminating the reliance on third-party encryption providers.
@@ IV-B MQFL-FHE: Multimodal Quantum Federated Learning with Fully Homomorphic Encryption
In MQFL-FHE, we employ an immediate feature-level fusion by utilizing multi-head attention mechanisms ((Vaswani, 2017)) to extract relevant features from diverse input modalities. The client-specific preprocessing function is represented as $\mathcal{D}_{k}\leftarrow\text{PrepareMultimodalDataset}(k)$, ensuring that varying input types are effectively managed and allowing local models to learn from different sources. The local model update for client $k$ can be expressed as:
@@ V Discussion
In the experimental setting of our study, we use datasets that vary significantly in terms of image or sequence count and imbalance factors to evaluate our models under a variety of conditions. The CIFAR-10 dataset, with a balanced distribution (imbalance factor of $1$), includes $60,000$ images and employs a batch size of $128$ for training, as shown in Table II. In contrast, the DNA Sequence dataset ((Singh, 2023)) presents a higher challenge with an imbalance factor of $5.595$ and a total of $4,380$ sequences, trained with a smaller batch size of $16$, to possibly mitigate the effects of its substantial class imbalance. Other datasets, such as MRI Scan and PCOS ((Hub et al., 2024)) with imbalance factors of $1.207$ and $2.544$, respectively, and total counts of $7,022$ and $3,200$, use a batch size of $32$. Regarding the hyperparameters, a learning rate of $1\times 10^{-3}$ is applied, which may increase to $3\times 10^{-3}$ if learning plateaus, as detailed in Table II. The FL utilizes the FLWR library ((Beutel et al., 2020)) to manage training across $10$ clients over $20$ communication rounds with $10$ epochs per client, contrasting with the $25$ total epochs in the centralized approach. Simultaneously, Pennylane ((Bergholm et al., 2018)) is employed to execute quantum computations within the same FL setup. The experiments are conducted on hardware with an AMD EPYC 7F72 (96) @ 3.685GHz CPU and NVIDIA A100-PCIE-40GB GPUs to ensure efficient processing and computation. This setup, incorporating diverse batch sizes and a range of hyperparameters, ensures comprehensive evaluation across both federated and centralized learning settings.
@@ VII-A Mathematical Analysis of Error Propagation and Noise Stabilization in FHE-Quantum Hybrid Models
Noise in FHE-based hybrid architectures arises primarily from errors introduced during the encryption and decryption processes of FHE. These errors accumulate as computations progress, and their propagation can significantly impact the fidelity of the results. Mathematically, such errors can be modeled as deviations in state evolution during computation, represented by transformations in the Hilbert space. However, the inherent properties of quantum operations, governed by the special unitary group $SU(2)$, impose strict constraints on the behavior of noise. Quantum operations on qubits correspond to rotations on the Bloch sphere, represented by the unitary matrix $U(\phi,\theta,\psi)$(Sultanow et al. [2024]). These operations adhere to the condition $U^{\dagger}U=I$, ensuring the preservation of the norm of the quantum state vector. Specifically, for a state vector $\vec{v}$, the transformation $\vec{v}^{\prime}=U\vec{v}$ satisfies:
@@
  • Error Correction Through Hybridization: The output of the quantum layer, with its stabilized noise, is passed to the classical neural network. The classical network processes the signal while suppressing residual quantum noise, leveraging the strengths of both paradigms to enhance overall error tolerance (Kundu and Ghosh [2024]).
@@ VII-B Effects of FHE Modifications
TABLE V: Impact of encryption parameters (bit scale, polymodulus degree, extrema count) on the number of encrypted values and computational time, highlighting trade-offs between security, efficiency, and scalability.
Bit Scale Polymodulus Degree Extrema Count Encrypted Parameter Count Time (s)
@@ VII-C Efficacy of MQMoE Approach
The dataset by Livingstone and Russo [2018] contains audio-visual recordings of 24 professional actors (12 male and 12 female), each expressing eight emotions: calm, happy, sad, angry, fearful, surprised, disgust, and neutral. The data is synchronized with both speech and song modalities, making it ideal for multimodal emotion recognition tasks that require the joint analysis of audio and visual cues. This richness in data supports the evaluation of the MQMoE approach in handling real-world, emotionally nuanced inputs across different modalities. As shown in Table VII, the findings from the RAVDESS dataset indicate that the MQMoE framework replicates trends observed in prior experiments. Specifically, the application of QFL+FHE combined yields enhanced performance metrics. The incorporation of audio-visual modalities improves recognition accuracy and proves the modularity and adaptability inherent in the MQMoE architecture. This demonstrates its effectiveness in multimodal emotion recognition, presenting a solution that can adapt to multiple modalities.
TABLE VII: Experiment results for RAVDESS Multimodal dataset across various setups.
Experiment Setup Test Loss Train Accuracy Test Accuracy Time (sec)
Classical Centralized $0.34$ $98.15\%$ $85.56\%$ $125.43\pm 0.08$
Quantum Centralized $0.373$ $96.82\%$ $87.28\%$ $459.10\pm 0.11$
Classical FL $0.74$ $97.95\%$ $76.58\%$ $670.57\pm 3.45$
QFL $0.93$ $97.30\%$ $70.83\%$ $1013.91\pm 1.81$
FL + FHE $1.16$ $95.55\%$ $67.43\%$ $892.14\pm 4.67$
QFL + FHE $0.83$ $94.53\%$ $76.43\%$ $1140.76\pm 1.69$

### v4

@@ I Introduction
Federated Learning (FL) has emerged as a powerful paradigm for collaboratively training machine learning models across decentralized data sources [1], allowing multiple clients to enhance model performance while preserving data privacy and complying with regulations such as the general data protection regulation (GDPR). FL alone is insufficient to fully protect data privacy, as the global model and updates shared between clients and the central server remain vulnerable to inference attacks. For instance, membership inference attacks use the model parameters or gradients to deduce whether a specific data sample was part of the training dataset [2]. Similarly, attribute inference attacks attempt to uncover sensitive attributes of individual data samples by exploiting correlations between model updates and data features [3]. Fully homomorphic encryption (FHE) was proposed by Gentry to mitigate vulnerabilities by allowing computations on encrypted data without decryption [4], ensuring that sensitive information remains secure even during model training and aggregation within new gen FL. It works well against gradient leakage attacks, in which hackers try to recover private information by looking at the gradients that are sent back and forth during training [5]. Adding FHE to FL adds a lot of extra work to computers, causing delays and less accurate models during the aggregation phase [6]. Consequently, while FHE protects data privacy, it can hinder the overall performance of FL systems, which is crucial for applications requiring high accuracy or real-time predictions. Despite the growing interest in multimodal FL, which aims to leverage various types of data from multiple clients, there exists a gap in research that addresses the unique challenges posed by FHE in this context [7]. Furthermore, the application of quantum computing (QC) to enhance FL frameworks has been largely unexplored. QC has the potential to generate richer representation within the federated system, which could alleviate performance bottlenecks introduced by FHE [8].
To mitigate the performance degradation that occurs during model aggregation, we propose a novel integration of QC with FHE in FL, taking into account the nature of the problem. Our method employs QC to counteract the model degradation resulting from FHE, allowing for an improved collection of model updates while protecting privacy, and for the first time, we present a novel multimodal quantum mixture of experts (MQMoE) model within the FL framework that employs FHE. We design this architecture to handle diverse multimodal data with MoE from various clients [9], enhancing representational and task-specific learning performance.
@@ II Related Works
QC in FL: Innan et al. introduced the federated quantum neural network (FedQNN) [10], which integrates quantum machine learning (QML) with QFL principles [11]. The experiments were conducted across genomics and healthcare datasets, laying the groundwork for investigations into quantum-enhanced learning methods in privacy-sensitive domains. This has been further advanced by efforts in QFL, which are creating new avenues in QML. In particular, multiple studies have emphasized using QFL in quantum systems and decentralized data frameworks [12, 13]. Applications in healthcare and other domains, using both quantum hardware and classical simulations, have shown promising results [14, 15, 16, 17, 18, 19]. Despite this progress, challenges related to resource allocation and efficient implementation persist [20, 21, 22], highlighting ongoing efforts to address these issues.
Privacy preservation in FL: Rahulamathavan et al. proposed FheFL [23], which incorporates a multi-key additive homomorphic encryption scheme to mitigate privacy risks associated with gradient sharing in FL. The experimental results confirmed that FheFL maintains model accuracy while providing robust security against potential data breaches, emphasizing the necessity of integrating FHE techniques into FL frameworks to enhance data privacy. Pan et al. introduced FedSHE [24], a useful FL scheme that uses adaptive segmented CKKS homomorphic encryption to keep gradients safe from leakage attacks. Their analysis of security parameters underscores the trade-offs between correctness, efficiency, and security in FL environments.
Multimodal data handling in FL: Yu et al. developed CreamFL [25], a framework designed for multimodal FL. This method enables clients with heterogeneous model architectures to contribute to a unified global model without sharing their private data. By implementing a contrastive representation ensemble strategy, CreamFL effectively addresses the challenges posed by modality discrepancies and task diversities. Gong et al. further advanced this area by introducing a multimodal vertical FL framework utilizing bivariate Taylor series expansion for encryption [26], thus eliminating the reliance on third-party encryption providers.
@@ IV-B MQFL-FHE: Multimodal Quantum Federated Learning with Fully Homomorphic Encryption
In MQFL-FHE, we employ an immediate feature-level fusion by utilizing multi-head attention mechanisms to extract relevant features from diverse input modalities [27]. The client-specific preprocessing function is represented as $\mathcal{D}_{k}\leftarrow\text{PrepareMultimodalDataset}(k)$, ensuring that varying input types are effectively managed and allowing local models to learn from different sources. The local model update for client $k$ can be expressed as:
@@ V Discussion
In the experimental setting of our study, we use datasets that vary significantly in terms of image or sequence count and imbalance factors to evaluate our models under a variety of conditions. The CIFAR-10 dataset, with a balanced distribution (imbalance factor of $1$), includes $60,000$ images and employs a batch size of $128$ for training, as shown in Table II. In contrast, the DNA Sequence dataset presents a higher challenge with an imbalance factor of $5.595$ and a total of $4,380$ sequences [28], trained with a smaller batch size of $16$, to possibly mitigate the effects of its substantial class imbalance. Other datasets, such as MRI Scan and PCOS with imbalance factors of $1.207$ and $2.544$ [29], respectively, and total counts of $7,022$ and $3,200$, use a batch size of $32$. Regarding the hyperparameters, a learning rate of $1\times 10^{-3}$ is applied, which may increase to $3\times 10^{-3}$ if learning plateaus, as detailed in Table II. The FL utilizes the FLWR library to manage training across $10$ clients over $20$ communication rounds with $10$ epochs per client [30], contrasting with the $25$ total epochs in the centralized approach. Simultaneously, Pennylane is employed to execute quantum computations within the same FL setup [31]. The experiments are conducted on a multi-GPU HPC node with an AMD EPYC 7F72 (96) @ 3.685GHz CPU to ensure efficient processing and computation. This setup, incorporating diverse batch sizes and a range of hyperparameters, ensures comprehensive evaluation across both federated and centralized learning settings.
@@ VII-A Mathematical Analysis of Error Propagation and Noise Stabilization in FHE-Quantum Hybrid Models
Noise in FHE-based hybrid architectures arises primarily from errors introduced during the encryption and decryption processes of FHE. These errors accumulate as computations progress, and their propagation can significantly impact the fidelity of the results. Mathematically, such errors can be modeled as deviations in state evolution during computation, represented by transformations in the Hilbert space. However, the inherent properties of quantum operations, governed by the special unitary group $SU(2)$, impose strict constraints on the behavior of noise. Quantum operations on qubits correspond to rotations on the Bloch sphere, represented by the unitary matrix $U(\phi,\theta,\psi)$ [32]. These operations adhere to the condition $U^{\dagger}U=I$, ensuring the preservation of the norm of the quantum state vector. Specifically, for a state vector $\vec{v}$, the transformation $\vec{v}^{\prime}=U\vec{v}$ satisfies:
@@
  • Error Correction Through Hybridization: The output of the quantum layer, with its stabilized noise, is passed to the classical neural network. The classical network processes the signal while suppressing residual quantum noise, leveraging the strengths of both paradigms to enhance overall error tolerance [33].
@@ VII-B Effects of FHE Modifications
TABLE V: Impact of encryption parameters on encrypted values and computational time, highlighting security, efficiency, and scalability trade-offs.
Bit Scale Poly. Deg. Extrema Enc. Count Time (s)
@@ VII-C Efficacy of MQMoE Approach
The dataset contains audio-visual recordings of 24 professional actors (12 male and 12 female) [34], each expressing eight emotions: calm, happy, sad, angry, fearful, surprised, disgust, and neutral. The data is synchronized with both speech and song modalities, making it ideal for multimodal emotion recognition tasks that require the joint analysis of audio and visual cues. This richness in data supports the evaluation of the MQMoE approach in handling real-world, emotionally nuanced inputs across different modalities.
@@
As shown in the results table, the findings from the RAVDESS dataset indicate that the MQMoE framework replicates trends observed in prior experiments. Specifically, the application of QFL+FHE combined yields enhanced performance metrics. The incorporation of audio-visual modalities improves recognition accuracy and proves the modularity and adaptability inherent in the MQMoE architecture. This demonstrates its effectiveness in multimodal emotion recognition, presenting a solution that can adapt to multiple modalities.